OpenStudy (anonymous):

Y- 1/3 = 2/9 how do I do this?

OpenStudy (anonymous):

Help me work it out please.

OpenStudy (anonymous):

y=2/9+1/3= 5 /9

OpenStudy (anonymous):

so y=5/9